Which equation represents a line whose x-intercept is the solution to the equation -3/4x+1/2=0
Answer Choices:
A) 3x+4y=2
B) 3x+4y=-1/2
C) 4x+3y=2
D) y=-3/4x-1/2

Answer by stanbon(48516)

-3/4x+1/2=0
Find the solution as follows:
(-3/4)x+(1/2)=0
Multiply thru by 4 to get:
-3x+2=0
x=2/3
---------
Determine which equation has an x-intercept of 2/3.
Check by letting y=0 and solving for x.
A) 3x+4y=2
3x+0=2
x=2/3
So, A is the answer.
Cheers,
Stan H.